### v1

# Explainable AI for Autism Diagnosis: Identifying Critical Brain Regions Using fMRI Data

Abstract
Early diagnosis and intervention for Autism Spectrum Disorder (ASD) has been shown to significantly improve the quality of life of autistic individuals. However, diagnostics methods for ASD rely on assessments based on clinical presentation that are prone to bias and can be challenging to arrive at an early diagnosis. There is a need for objective biomarkers of ASD which can help improve diagnostic accuracy. Deep learning (DL) has achieved outstanding performance in diagnosing diseases and conditions from medical imaging data. Extensive research has been conducted on creating models that classify ASD using resting-state functional Magnetic Resonance Imaging (fMRI) data. However, existing models lack interpretability. This research aims to improve the accuracy and interpretability of ASD diagnosis by creating a DL model that can not only accurately classify ASD but also provide explainable insights into its working. The dataset used is a preprocessed version of the Autism Brain Imaging Data Exchange (ABIDE) with 884 samples. Our findings show a model that can accurately classify ASD and highlight critical brain regions differing between ASD and typical controls, with potential implications for early diagnosis and understanding of the neural basis of ASD. These findings are validated by studies in the literature that use different datasets and modalities, confirming that the model actually learned characteristics of ASD and not just the dataset. This study advances the field of explainable AI in medical imaging by providing a robust and interpretable model, thereby contributing to a future with objective and reliable ASD diagnostics.
Index Terms:
Biomarkers, Deep Learning, Interpretability, Sparse Autoencoders, RemOve And Retrain (ROAR)
I Introduction
Autism Spectrum Disorder (ASD) is a complex neurodevelopmental condition characterised by impairments in social communication, interaction, and restricted, repetitive behaviors [1]. Its presentation varies widely across affected individuals, impacting on quality of life. In the last 20 years, there has been a 787% surge in ASD diagnosis increasing its prevalence from approximately 1% to around 3% of the population. This surge in diagnosis is thought to be due to the delay in recognition of the clinical symptoms and the lack of personnel to diagnose. [2, 3, 4]. Early diagnosis, ideally between the ages of two and five, can lead to a significantly better Quality of Life (QoL) for individuals with ASD [5]. Studies also show that early intervention can improve IQ scores by up to 18 points and enhance language skills [6], emphasising the critical need for early diagnosis and intervention. However, the average age of diagnosis remains around five years old, highlighting a crucial limitation in current diagnostic practices [7].
Specifically, current clinical ASD diagnostic practices lack effective early diagnosis methods. The method of diagnosing ASD in children is behavioral analysis combined with clinical history, which is not effective for early diagnosis as it is based on identifying abnormal communication and social behaviors that may not emerge until the condition is well established [8]. It is also subject to clinicians’ biases and has historically led to under-diagnosing ASD particularly, in certain groups such as females [9]. To enhance our understanding of ASD and improve diagnostic accuracy, there is a need for objective, biologically-based markers of the condition.
The resting-state functional Magnetic Resonance Imaging (fMRI) data of people with ASD has been analysed to show abnormalities in functional connectivity [^1] Functional connectivity examines the correlation of activity between different brain regions. If two areas consistently show increases and decreases in activity together, they are considered functionally connected. [10]. These differences in distinctive patterns in the functional connectivity of autistic individuals can serve as biomarkers of ASD. By researching these differences, studies have identified several brain networks and areas that are associated with ASD such as disrupted connectivity in the Default Mode Network (DMN) [11]. However, the vast differences in ASD and the concept of the “idiosyncratic brain” suggests that there exists a highly personalised neural network configuration for each individual [12, 10]. The regions involved are also affected by the severity of ASD [13]. This highlights the need for a method to identify the intricate patterns of neural networks in ASD and possibly reveal a biomarker that is characteristic of all ASD individuals regardless of the severity of ASD.
These abnormalities can be learned using deep learning (DL) to distinguish between ASD and Typical Control (TC). Many recent studies have emerged using functional connectivity for ASD classification using machine learning, with some showing a high accuracy ($>$ 90%) [14, 15]. While these models cannot replace clinical professionals, they can offer valuable decision support. These neuroimaging-based diagnosis methods have the potential to revolutionise ASD diagnosis, but most current approaches lack in explainability. These limitations reduce trust and hinder clinical acceptance of the model in real-world scenarios [16]. To address this challenge, there is a large body of research focusing on approaches that enhance the interpretability of DL models in the medical domain [17]. However, identifying which approach should be used with functional connectivity data or how they should specifically be used to identify biomarkers in ASD has not yet been explored.
This study explores the utilisation of DL methods to identify potential biomarkers for ASD through the analysis of resting-state fMRI data, aiming to develop and explain a predictive model that is both accurate and interpretable. We also benchmark a variety of interpretability methods using the Remove And Retrain (ROAR) technique [18] to identify the most effective method for finding important regions in the brain for ASD classification.
This research has the potential to significantly improve our understanding of the neurobiological basis of ASD. This study also contributes to the growing field of explainable AI in healthcare, addressing the need for transparency and interpretability in medical AI systems, as required by regulatory bodies such as the FDA [19]. Our findings are critically evaluated in the context of existing research and potential limitations are discussed. A comprehensive analysis of our model’s performance and its implications for ASD research and diagnosis is provided. The regions highlighted by our model are compared with existing literature to verify whether our model genuinely learned about ASD characteristics or simply learned patterns specific to our dataset.
Successful application of the outcomes of this research could lead to improved access to early interventions, reduced costs of ASD diagnosis, and the development of more effective diagnostic tools. By enhancing our understanding of the neurobiological markers of ASD, this work aims to contribute to the development of more targeted interventions and improved quality of life for individuals with ASD and their families.
II Related Work
II-A Neurological basis of ASD and fMRI studies
Research into ASD using fMRI has revealed distinct connectivity patterns between individuals with ASD and typical controls. Understanding these patterns is vital as they provide objective, neurobiological markers that could enhance early detection and inform targeted interventions. Functional connectivity, which measures synchronised activity between different brain regions, shows atypicality in several areas frequently linked to ASD. Key regions include the frontal cortex, involved in executive functions and social cognition [20], temporal regions for language and social processing [21], the amygdala for emotion processing [22], and the cerebellum for motor function and social cognition [23]. Studies indicate that ASD individuals often exhibit local over-connectivity and long-range under-connectivity [24, 25, 26]. Furthermore, disrupted connectivity in the DMN and Salience Network (SN) is associated with ASD [11, 27]. However, there is considerable variation in the important regions identified across studies, and these regions can change on a case-by-case basis depending on the severity of ASD symptoms [13]. Notably, a universal biomarker common to all cases of Autism, if it exists, has not been found. The concept of the “idiosyncratic brain” suggests highly personalised neural network configurations in ASD, presenting a complex challenge for traditional analysis methods. This emphasises the need for advanced techniques like DL [^2] Traditional ML methods struggle with the high dimensionality and complex, non-linear relationships present in neuroimaging data. Deep learning, with its ability to automatically learn hierarchical representations from raw data, is particularly well-suited for capturing the subtle and highly variable connectivity patterns observed in ASD. to identify the intricate patterns in ASD neuroimaging data [12].
II-B Deep Learning for brain imaging
DL techniques, inspired by the structure of the brain, have proven highly effective in medical imaging, surpassing traditional machine learning methods [28]. DL models capture complex patterns and combine data from multiple modalities, improving diagnostic accuracy. In brain imaging, various DL architectures like Convolutional Neural Networks (CNNs), Recurrent Neural Networks (RNNs), Graph Neural Networks (GNNs), and Autoencoders are employed for different analysis tasks [29, 30, 31]. Challenges include managing high-dimensional and multimodal data, mitigating overfitting, and ensuring model interpretability. Initiatives promoting data sharing, ethical frameworks, and cloud-based resources are addressing these challenges, while explainable AI research is gaining traction to identify key brain regions driving DL model predictions. Ethical considerations are paramount, as DL models must be evaluated rigorously to prevent biases and discriminatory practices [32]. In the context of ASD diagnosis, these advanced DL techniques show great promise, offering the potential for high accuracy and early detection by leveraging complex patterns in brain imaging data that may not be apparent to the human eye or traditional analysis methods.
II-C Feature analysis and interpretability in ASD research
Feature analysis helps translate a model’s accuracy into insights about the underlying mechanisms it uses for classification. Methods include gradient-based techniques like Layer-wise Relevance Propagation (LRP) [33] and DeepLift [34], and perturbation techniques like SHapley Additive exPlanations (SHAP)[35] and Local Interpretable Model-Agnostic Explanations (LIME) [36]. In medical research, these interpretability methods have been successfully applied to identify biomarkers for various conditions [37, 38, 39]. However, a comprehensive review [40] found that there is no universally “best” interpretability method; rather, their effectiveness varies depending on the specific application and data type.
In the context of ASD research, there is a notable gap in the application of these interpretability methods, particularly for functional connectivity data derived from fMRI. This gap is exacerbated by several challenges unique to ASD and neuroimaging data. Firstly, the heterogeneity of ASD manifestations adds another layer of complexity, as different neural patterns may be significant for different subgroups within the ASD spectrum. Secondly, the lack of ground truth for important brain regions in ASD makes it challenging to validate the results of interpretability methods. Moreover, even when studies attempt to incorporate interpretability, they often fail to validate their results, leaving uncertainty about whether the model is truly learning characteristics of Autism or merely recognising patterns specific to the dataset. While some studies have attempted to use similar approaches in related fields, such as the use of back-propagation to identify biomarkers for Attention Deficit Hyperactivity Disorder (ADHD) [41], these efforts often suffer from limitations like low accuracy and lack of proper benchmarking. The absence of a standardised framework for evaluating and comparing different interpretability methods in the context of ASD further complicates the situation. These challenges collectively highlight the critical need for a systematic approach to evaluate and apply interpretability methods in ASD research, which is precisely what this study aims to address.
II-D Gaps in current research
A significant challenge in ASD research is the lack of reliable biomarkers, which hinders early diagnosis. Despite numerous studies, a universal biomarker common to all cases of Autism has not been identified, largely due to the heterogeneous nature of ASD and the notion of unique neurological variations among individuals [12]. This absence of definitive neurobiological markers significantly impedes the development of objective, early diagnostic tools for ASD.
Many studies in ASD diagnosis using DL face methodological challenges. Notably, there’s often a limited focus on evaluation metrics, with many studies reporting only accuracy, ignoring other crucial metrics like sensitivity and specificity [42]. This narrow focus can lead to overfitting and a false sense of model performance. Moreover, while some studies have achieved high accuracy, there’s a lack of explainable models with accuracy exceeding 90%. This gap raises a critical question: how can we trust an explainable model if it’s not sufficiently accurate, and conversely, how can we rely on a highly accurate model if we can’t explain its decisions?
The majority of studies have not employed interpretability methods to elucidate the regions utilised by their models. This results in “black box” models that may achieve high performance but lack a mechanism to verify whether they are genuinely learning to identify Autism or merely memorising dataset-specific patterns. Furthermore, when interpretability methods are used, they are often not benchmarked against each other, leaving uncertainty about which methods are most reliable for ASD research.
Another critical issue is the lack of diverse datasets representing various ethnicities, socioeconomic backgrounds, and ASD severity levels, which limits the models’ ability to generalise across diverse populations [43]. This could lead to biased diagnoses and exacerbate existing health inequalities if implemented in clinical settings without proper validation.
This project aims to address these critical gaps by developing an explainable DL model for ASD classification with high accuracy, rigorously benchmarking interpretability methods, and identifying key brain regions contributing to the diagnosis. While addressing the diversity issue is crucial for future clinical applications, it is beyond the scope of this research due to limited resources and the constraints of available datasets. Nevertheless, our study paves the way for more reliable, transparent, and potentially generalisable AI-assisted ASD diagnostics.
III Methodology
III-A Data Collection and Preprocessing
III-A1 Dataset
This study used the Autism Brain Imaging Data Exchange (ABIDE) dataset [44], specifically the Preprocessed Connectomes Project (PCP) version. ABIDE was chosen for its large sample size, ease of access, and prevalence in comparable studies, facilitating performance benchmarking.
The dataset includes resting-state fMRI data from 539 individuals with ASD and 573 typically developing controls, aged 7-64 years. While ABIDE offers a robust sample size for model training, it has a notable gender bias (85.2% male) [45]. Notably, We employed stratified k-fold cross-validation during model training and evaluation, ensuring that each fold maintained the original dataset’s distribution. We acknowledge that our results should be interpreted with caution regarding generalisability across genders. The data is anonymised in accordance with HIPAA guidelines and 1000 FCP / INDI protocols, ensuring ethical use of the neuroimaging data.
III-A2 Preprocessing
The PCP openly shares preprocessed neuroimaging data from ABIDE. There is no set standard on how neuroimaging data should be preprocessed so PCP offers data preprocessed using four different preprocessing pipelines, namely the Connectome Computation System (CCS) [46], the Configurable Pipeline for the Analysis of Connectomes (CPAC) [47], the Data Processing Assistant for Resting-State fMRI (DPARSF) [48] and the NeuroImaging Analysis Kit (NIAK) [49]. The preprocessing steps carried out by the various pipelines are comparatively analogous [14]. The main differences are in the particular algorithms for each step, the software simulations, and the parameters applied. Details of each pipeline technique are provided on the PCP website [50]. Since there is no consensus on what precprocessing pipeline is the best we have trained a model using each of the four pipelines and compared their results.
PCP applies some standard low-level fMRI preprocessing steps to all preprocessing pipelines. These include brain extraction, spatial smoothing, temporal filtering, realignment (motion correction), slice timing correction and intensity normalisation [47]. Further, there has been significant discussion in the literature about whether Framewise Displacement (FD)  [^3] Framewise Displacement (FD) is a measure of how much head movement occurs during the MRI scan filtering affects the accuracy of an autism diagnosis model. This necessitates an investigation to clarify head movement’s role in model performance. Some influential studies state that an FD value $>$ 0.2mm corrupts the fMRI data [51, 52], leading to a decrease in performance. In contrast, another study stated that FD filtering or head movement did not affect the performance of their model [53]. By conducting our study under both conditions - with and without FD filtering - we aim to provide clarity on this methodological issue and its impact on model accuracy.
III-A3 Parcellation
Brain images are overlaid on a brain atlas, which divides the brain into several Regions of Interest (ROIs). The voxels within each ROI are averaged to produce a single value, effectively reducing the dimensionality of the data. For this project, we selected the Automated Anatomical Labelling (AAL) atlas, which offers a balanced approach with 116 ROIs. This choice was driven by computational efficiency and proven performance. The computational cost of our model increases quadratically with the number of ROIs due to the use of functional connectivity matrices (FCM, explained in III.B), and the AAL atlas provides a good balance between granularity and manageable computational demands. Moreover, the AAL atlas has demonstrated high accuracy in similar studies [54, 15], supporting its suitability for ASD classification tasks. After preprocessing and atlas application, our dataset comprises of 884 samples (408 ASD, 476 TC). Each sample is represented as a 2-dimensional array $(N\texttimes t)$, where $t$ is the temporal dimension and $N$ corresponds to the number of ROIs in our atlas (116 in our case).
III-B Feature Extraction
To reveal the patterns of coordination between brain regions, the data is converted to a functional connectivity matrix (FCM) using Pearson’s correlation coefficient $R(i,j)$ between the intensity time series from two regions of interests (ROIs). Furthermore, Fisher’s Z transformation  [^4] Fisher’s Z transformation is a statistical method used to convert Pearson’s correlation coefficient to a normally distributed variable. This transformation is particularly useful when working with correlation coefficients, as it stabilises their variance and makes them more suitable for statistical analyses. was applied to the data.
The Fisher’s z-transformed correlation coefficients were then represented as a 116 x 116 matrix which was symmetric with respect to the diagonal due to the symmetric nature of correlation matrices $R(i,j)=R(j,i)$. To reduce redundancy, the lower triangular matrix of the FCM was taken and flattened to form a feature vector for each of the $884$ samples creating a feature vector with $6670$ features.
Creating a feature vector from the FCM has been thoroughly researched and used in DL models, however, when using them directly for autism diagnosis, the DL models peaked at an accuracy of 66.8% [55, 56]. This implies that these DL models suffer from the high-dimensionality and low number of samples problem. To efficiently use DL models, feature dimensionality must be reduced, or sample size must be increased. However, increasing sample size is infeasible due to reasons including high MRI costs and difficulty in participant recruitment.
Support Vector Machine with Recursive Feature Elimination (SVM-RFE) was used to reduce the feature set from $6670$ features to a 1000 features per sample [15]. This choice considers computational constraints, proven performance and the fact that SVM-RFE is a supervised feature selection approach, whereas unsupervised methods like Principal Component Analysis (PCA), operate without knowledge of the classification task, which can affect the accuracy.
III-C Deep Learning Model
This study utilises a DL architecture (Figure 1) that comprises of a Stacked Sparse Autoencoder (SSAE) combined with a softmax classifier to classify ASD using functional connectivity data.
The SSAE was chosen for this study due to its ability to effectively handle high-dimensional functional connectivity data. Autoencoders are neural networks that learn to compress input data into a lower-dimensional representation and then reconstruct it, making them ideal for capturing essential features in complex datasets. The “sparse” aspect is especially important for ASD classification, as it encourages the model to focus on the most distinctive biomarkers within the functional connectivity data while reducing noise and overfitting. The hierarchial structure of the SSAE, consisting of multiple layers of sparse autoencoders, encourages the model to learn increasingly abstract features of the data with each layer. This is important since biomarkers in functional connectivity may involve complex multi-level patterns.
To address potential overfitting and improve generalisability, we incorporate a sparsity constraint using the Kullback-Leibler (KL) divergence  [^5] Kullback-Leibler (KL) divergence is a measure from information theory that quantifies how different one probability distribution is from another term in the loss function:
$L(x,\hat{x})=MSE+\beta\times KL(p\|\hat{\rho_{j}})$ (1)
where $MSE$ is the mean squared error between input $x$ and its reconstruction $\hat{x}$, $\rho$ is the desired sparsity, $\hat{\rho}$ is the average activation of a neuron, and $\beta$ controls the weight of the sparsity penalty.
Figure 1: The overall architecture of the proposed approach, incorporating both the SSAEs and softmax classifier.
Our SSAE consisted of two autoencoder layers that used the ReLU activation function. The first reduces 1000 input features to 500, and the second further compresses to 100 neurons. The output was then fed into a softmax classifier that was trained using cross-entropy loss for ASD/TC classification. This is visualised in Figure 1.
III-D Training and Evaluation
The model was trained in two stages combining the strengths of both supervised learning and unsupervised learning. To ensure robust evaluation and training a stratified k-fold cross-validation approach with $k$ value of 5 was employed. A k-fold value of 5 was chosen to balance computational cost with variance reduction in performance estimates with regard to the dataset size of 884. A higher $k$ value led to there not being enough test cases for the model to be tested properly and a lower $k$ value leads to there being not enough training data and the model struggles to generalise. Within each fold, the dataset was split into training, validation, and testing sets, with careful consideration given to maintaining a balance between computational efficiency and robust performance assessment.
III-D1 Unsupervised Pre-training
The SSAE underwent unsupervised pre-training using the greedy layer-wise training approach [57]. This sequential approach allows each autoencoder layer to learn increasingly abstract representations of the input data. SAE 1 was trained first and its compressed hidden layer representation then served as input for training SAE 2. Both layers were trained using the Adam optimiser, chosen for its adaptive learning rate and faster convergence [58]. Performance monitoring during pre-training combined Mean Squared Error (MSE) and a sparsity penalty achieved through KL Divergence. This approach has shown efficacy in ASD classification, potentially improving generalisation and reducing overfitting [56, 59, 15].
III-D2 Supervised Fine-Tuning
Following pre-training, the SSAE’s encoder layers were combined with a softmax classifier for supervised fine-tuning. This stage optimises classification performance by backpropagating the error signal [^6] The error signal is the cross entropy loss from the softmax classifier, not the reconstruction loss used to train the encoder layers in the pre-training step. from the softmax classifier, modifying the encoder weights to improve discrimination between ASD and TC classes. Unlike many previous approaches that focused solely on fine-tuning the classifier, our study fine-tunes both the encoder layers and the classifier, which may contribute to higher classification accuracy [55, 56, 59].
The fine-tuning stage leveraged a hybrid training approach, combining the SSAE’s unsupervised representation learning capabilities with supervised training for the specific ASD classification task. This process employed the Adam optimiser [^7] Hyperparameters: Pre-training - 50 epochs, learning rate of 0.001, weight decay of 0.0001, batch size of 128, KL Divergence parameters: $\rho$ = 0.2, $\beta$ = 2. Fine-tuning - 50 epochs, learning rate of 0.0001, weight decay of 0.0001, batch size of 128., chosen for its demonstrated efficiency in the pre-training phase. The hyperparameters were carefully selected to balance performance gains against potential overfitting, allowing for subtle adjustments to the pre-trained weights while maintaining regularisation.
III-D3 Performance Evaluation
A rigorous evaluation of the models’s performance on the ASD classification task is necessary to assess its effectiveness and generalisability. We used the previously established stratified k-folds and splits to maintain consistency. Key metrics, including accuracy, sensitivity, specificity, precision, and F1-score, were calculated for each fold and then averaged to provide an overall assessment of the model’s classification ability. Additionally, loss graphs were visualised for training and validation to avoid and monitor overfitting. This evaluation seeks to determine whether the model can surpass the performance limitations often observed in similar ASD classification models while avoiding overfitting.
III-E Feature Analysis and Interpretation
This section explores the different methods that were used to extract meaningful features used by the model for its diagnosis. Understanding the basis of its decisions is essential for trust and clinical translation.
III-E1 Interpretability Methods
To identify potential neurobiological markers of ASD and enhance the model’s explainability, a rigorous feature analysis was conducted. This investigation involved multiple interpretability techniques, including:
  • Integrated Gradients [38]: Calculates feature importance by examining how changes in input values impact the predictions
  • LIME [36]: Approximates complex model decisions locally through perturbations of the input data.
  • SHAP [35]: A game-theoretic approach attributing feature importance based on its contribution to the model’s output.
  • DeepLift [34]: Decomposes the output prediction by backpropagating the contributions of neurons against a reference point.
  • DeepLiftShap [60]: Combines DeepLift’s insights with SHAP values for enhanced explainability.
  • GradientShap [35]: Computes SHAP values using gradients, offering computational advantages.
  • GuidedBackprop [61]: Highlights input features relevant to the prediction by visualising positive gradients during backpropagation.
These methods were chosen as they use diverse approaches to enhance explainability. Gradient-based techniques (Integrated Gradients, DeepLift, DeepLiftSHAP, GradientSHAP) directly examine feature influence on predictions. The LIME approach explores model sensitivity to input perturbations, while the SHAP approach employs a game-theoretic approach for feature attribution. Methods like DeepLiftSHAP and GuidedBackprop combine strengths for greater robustness. As each method may yield varying feature rankings due to their differing focus, a benchmark is crucial for trustworthy interpretation.
III-E2 Remove and Retrain (ROAR)
To systematically compare these methods, the Remove and Retrain (ROAR) benchmark was employed [18]. ROAR works by:
1. Identifying Important Features: Each interpretability method generates a ranking of input features, showing which features the method considers most important for making predictions.
2. Testing Importance: ROAR tests how crucial those top features truly are. It takes the top ’t’ percent of features (using different threshold values like 20%, 40%, etc.) and replaces them with meaningless values.
3. Retraining and Comparing: For each threshold, the model is retrained on the modified data, and its performance is measured. If an interpretability method is reliable, removing its top-ranked features should cause a steep decline in the model’s performance compared to removing features identified by other methods. Once these are computed they are displayed as a line graph to see which method had the steepest decline and that method is noted as the best performing method. Additionally, the results are compared to a baseline where features are removed randomly. This ensures that any observed performance drops are due to the interpretability method identifying truly important features, and not just the general effect of removing any features.
For each interpretability method the top features were ranked, then during the ROAR analysis the top $t$ percent of features were replaced by zero. Zero was chosen as the uninformative value because the input features are correlation coefficients and zero correlation indicates that no positive or negative relationship exists between the two variables. Hence, zero is uninformative in the context of functional connectivity. The threshold values of $t$ were chosen to focus on studying the importance of the top 10% of features in ASD classification while still monitoring the rest of the features  [^8] The threshold values of $t$ chosen for this research are: $[0.01,0.05,0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,0.95,0.99]$. The results are analysed in section IV.
III-E3 Important Regions of Interest (ROIs)
Figure 2: Important ROIs (yellow) identified by the Integrated Gradients approach when using CPAC pipeline.
Figure 3: Important ROIs (yellow) identified by the Integrated Gradients approach when using CCS pipeline.
Figure 4: Important ROIs (yellow) identified by the Integrated Gradients approach when using DPARSF pipeline.
Once ROAR established the reliability of some interpretability methods, the focus was moved to the brain regions (ROIs) represented by the features identified by the method to understand their role in ASD.
Each feature of the model is a correlation coefficient which means each feature is a connection between two ROIs. To calculate individual important ROIs, weights were calculated by multiplying the frequency of that ROI in the top connections and its highest weight. These individual weights were also graphed with a larger size and brighter color of node indicating more importance showcased in Figures 4, 4 and 4. Once the most important ROIs for each preprocessing pipeline were determined, each region was mapped to its Brodmann Area and the overlap between the different preprocessing pipelines were studied.
IV Results and Discussion
The model achieved state-of-the-art performance on the ABIDE dataset, with a maximum accuracy of 98.2% across three different preprocessing pipelines, surpassing previous models for ASD classification (Table II). Evaluation metrics, including specificity, precision, and F1-score, were high with values of 0.98, 0.98 and 0.97 respectively. Analysis of the model’s decision-making process identified a set of discriminative functional connectivity features. These results are presented in detail, alongside a discussion of model performance, interpretability performance, feature analysis and the potential significance of the identified ROIs.
IV-A Model performance
Figure 5: Loss graphs for our model in the first K-Fold. AE1, AE2 and classifier loss is reconstruction loss during unsupervised pre-training. Model loss and accuracy is during supervised fine-tuning.
Figure 6: Performance of the model across different preprocessing pipelines averaged across 5 K-folds.
Model Accuracy
Model without FD Filtering 91%
Model with FD Filtering 98.2%
TABLE I: Model accuracy with and without framewise displacement filtering of samples $>$ 0.2mm FD.
The model demonstrated exceptional performance on the ASD classification task, particularly when samples with a FD value $>$ of 0.2mm were filtered. The 7% decrease in accuracy (Table I) along with the observation that the autoencoders could not accurately reconstruct the data when a sample has more than 0.2mm FD, suggests that the hypothesis regarding FD’s impact on data quality is indeed valid. This finding aligns with previous research on the subject [52] [51]. In particular, all the incorrect predictions during testing also had a high FD, suggesting that head movement during MRI has a significant effect on the precision of the diagnosis of ASD. This finding contrasts with previously reported results on the subject [53].
Using FD Filtering, an accuracy of 98.2% using the DPARSF pipeline (averaged across 5 K-Folds) was achieved by the model. These metrics surpass all previous state-of-the-art models (Table II) using the fMRI modality, reported as 93.59% accuracy in [15]
Notably, CCS, DPARSF and CPAC all performed within 1% of each other while NIAK trailed behind in every metric by an average of 4% (Figure 6). This difference in performance likely stems from the different approaches used for NIAK compared to CCS, CPAC, and DPARSF. In the NIAK pipeline, the functional data was written into template space prior to calculating the statistical derivatives, whereas for the others the statistical derivatives were calculated in native space and then the functional data was written into template space [50].
Model Accuracy(%)
Pavithra et al. (2023) [62] 85.0
Bhandage et al. (2023) [63] 92.4
Wadhere et al. (2023) [64] 88.1
Wang et al. (2019) [15] 93.5
Our model 98.2
TABLE II: Accuracy comparison with papers that have a dataset $>$ 500 samples using the fMRI modality.
The loss graphs (Figure 5) reveal an interesting distinction. During the unsupervised pre-training stage, the loss graphs of AE 1 and AE 2 show similar losses during training and validation. Since this is reconstruction loss and the AEs can accurately reconstruct the data, we can infer that the dataset is mostly homogenous and the distinctions between samples are minuscule when not differentiating between ASD and TC which is expected since the difference between a brain with ASD and one without is subtle [65]. During fine-tuning, we use a cross-entropy loss which results in a more typical graph where the validation loss is slightly higher than the training loss, but both are decreasing. This pattern suggests that the model is learning continuously without overfitting, which is further supported by the accuracy graph.
The model’s success in classifying ASD likely stems from its hybrid architecture. The combination of unsupervised pre-training and supervised fine-tuning appears to have enabled the model to learn meaningful representations of complex connectivity patterns. This aligns with findings suggesting that DL models benefit from such multi-stage training approaches, especially in medical image analysis [66].
IV-B Feature Importance
IV-B1 ROAR Benchmark
The ROAR analysis identified which interpretability methods are reliable. It showed that Integrated Gradients, Deep fLift, DeepLiftShap, and GradientShap methods all had similar results and showed the steepest drop in accuracy when the initial features were removed, indicating that the gradient-based methods identified the most discriminative features required for ASD classification (Figure 7). The Integrated Gradients approach showed the steepest drop in accuracy and was therefore assumed the most reliable method. Notably, the LIME approach did not perform much better than the random baseline, indicating that feature perturbation-based techniques do not perform well in the functional connectivity field. This contrasts with results found in the original ROAR study [18] that showed that the Integrated Gradients approach performed similarly to the random baseline, further emphasising that the best-performing interpretability methods change depending on the model architecture and the data’s modality.
Figure 7: ROAR Analysis, shows how removal of top 1% of features leads to steeper accuracy drop for the Integrated Gradients, DeepLift, DeepLiftShap and GradientShap methods.
IV-B2 Feature Analysis
Feature analysis using the Integrated Gradients approach revealed a set of consistent brain connectivity patterns highly discriminative for the classification of ASD across the different preprocessing pipelines which can be observed in Figures 4, 4 and 4.
Important brain areas (parcellated per the AAL atlas), identified by the interpretation method are shown in columns CCS, CPAC, and DPARSF in Table III. The regions are arranged in descending order of importance score, with their corresponding Brodmann Areas [67]. The most consistently important regions across pipelines are areas 17 (Primary visual cortex) and 18 (Secondary visual cortex), with areas 41 and 42 (Auditory cortex) also showing significance.
The consensus results across the three pipelines suggest that ROIs involved in visual processing are central to the classification of participants as autistic versus controls. The occipital lobe identifies three visual regions of particular importance: the calcarine sulcus, cuneus, and superior occipital gyrus.
Notably, all three pipelines rank the calcarine sulcus as the most important area. This finding is particularly significant as it aligns with current evidence from multiple research dimensions, including genetic studies that have independently highlighted the importance of Brodmann areas 17 and 18 in autism [68]. Other research dimensions including studies on global motion perception deficits [69] and atypical gamma oscillations [70] show ASD being associated to the primary visual cortex. Our model’s identification of this region corroborates growing evidence that fundamental differences in early visual processing may play a crucial role in ASD, potentially influencing broader social and behavioral symptoms.
Similarly, the cuneus (part of Brodmann area 17) emerges as a significant region. Recent research has shown that autistic individuals have reduced connectivity between the brainstem and the cuneus compared to their non-autistic co-twins, potentially indicating changes in the low-level visual pathway that may affect social information processing and low-level perception [71]. Furthermore, fMRI and eye-tracking studies have implicated the cuneus in social dysfunction in autism [72].
The pipelines also agree on the importance of the angular gyrus, located in the posterior part of the inferior parietal lobe, which is involved in language processing. This finding aligns with the known language and communication difficulties often observed in ASD [73].
The convergence of our machine learning-based findings with these different studies provides strong validation that our model has indeed captured genuine neurobiological markers of ASD, rather than dataset-specific patterns. These results also demonstrate its potential to corroborate and extend existing knowledge in autism research. The consistency of our findings across different preprocessing pipelines and their alignment with diverse research approaches (genetic, neuroanatomical, and functional) highlights the robustness of our method and its potential for advancing our understanding of the neurobiological basis of ASD. Figure 8 highlights these important regions, providing a visual representation of the key areas identified by our analysis.
Importantly, while these top ROIs offer potential insight, it is crucial to remember that ASD is a complex disorder. Significant patterns likely arise from a combination of interacting brain regions. Further investigation might focus on broader network-level differences rather than isolated ROIs. It’s noteworthy that our dataset included a variety of ASD severity cases, and the identified biomarkers were common among all of them. This suggests that these visual processing regions may represent a fundamental neurobiological signature of ASD, potentially present across the spectrum of severity.
CCS Brodmann Area CPAC Brodmann Area DPARSF Brodmann Area
Calcarine_L 17 Calcarine_R 17 Calcarine_L 17
Occipital_Sup_R 17 Cuneus_R 17,18 Rectus_R 11,12
Occipital_Mid_R 17 Occipital_Sup_R 17 Cuneus_R 17,18
Cuneus_R 17,18 Frontal_Inf_Tri_R 44,45 Temporal_Mid_L 21
Angular_R 39 Angular_L 39 Cerebelum_Crus1_R -
Cerebelum_8_L - Cerebelum_8_R - Calcarine_R 17
Thalamus_R 41,42 Cerebelum_Crus2_L - Cerebelum_3_R -
Calcarine_R 17 Calcarine_L 17 Temporal_Inf_L 20
Insula_R 13,16 Lingual_L 17 Occipital_Sup_R 17
SupraMarginal_R 40 ParaHippocampal_L 34 Frontal_Sup_Medial_R 11,12
Parietal_Inf_L 5 Precentral_R 4 Frontal_Sup_L 11,12
Heschl_R 41 Thalamus_L 41,42 Angular_L 39
TABLE III: The top ROIs identified by each pipeline (CCS, CPAC and DPARSF) in decreasing order of importance alongside its corresponding Brodmann Area.
Figure 8: Shows the Brodmann’s area[74]. The different colors represent different brain functions. The common important brain regions from the table III are highlighted.
V Conclusion
This research significantly advanced the field of ASD classification using deep learning models applied to resting-state fMRI data, addressing several key gaps in current literature. Our model achieved exceptional performance with an average accuracy of 98.2% using the DPARSF pipeline, surpassing the previous benchmark of 93.59% [15]. This addresses the critical gap of lacking an explainable model with over 95% accuracy in ASD classification. The model’s success stems from its hybrid architecture, combining unsupervised pre-training and supervised fine-tuning, and the removal of corrupt data. This helped the model effectively capture complex functional connectivity patterns characteristic of ASD.
Unlike many previous studies that focused solely on accuracy, we provided comprehensive evaluation metrics including sensitivity, specificity, precision, and F1 scores. This thorough approach offers a more robust assessment of the model’s performance and addresses the limitation of narrow evaluation practices in the field. Our study also revealed the significant effect of head movement on the precision of ASD diagnosis, contrasting with some previous findings and highlighting the importance of motion correction in fMRI studies of ASD.
We pioneered the use of ROAR (Remove And Retrain) to benchmark different deep learning interpretability methods in the context of functional connectivity data. This novel application revealed that gradient-based methods, particularly Integrated Gradients, are most suitable for this type of data, filling a significant gap in understanding the reliability of various interpretability approaches in ASD research.
Our feature analysis highlighted the importance of visual processing regions, particularly the calcarine sulcus and cuneus, in ASD classification. These findings were consistent across different preprocessing pipelines and aligned with independent genetic studies [68], providing strong validation of our model’s ability to capture genuine neurobiological markers of ASD rather than dataset-specific patterns. Notably, our identified biomarkers were common among a variety of ASD severity cases in our dataset, suggesting these visual processing regions may represent a fundamental neurobiological signature of ASD across the spectrum of severity.
In summary, this research has made significant strides in the application of deep learning to ASD classification and biomarker identification. By developing an accurate, explainable model and rigorously benchmarking interpretability methods, we have not only advanced the technical aspects of ASD diagnosis but also contributed to the neurobiological understanding of the disorder. The consistent identification of visual processing regions as key discriminators across different severity levels of ASD opens new avenues for research and potentially for clinical applications. While challenges remain, this work represents a substantial step forward in the quest for objective, reliable, and interpretable methods for ASD diagnosis and characterisation.
VI Limitations and Future Directions
Despite the promising results of this study, several limitations must be acknowledged, each of which points to important directions for future research. A primary limitation is the binary nature of our classification model. By focusing on distinguishing between ASD and TC and due to a dataset limitation, we may have inadvertently created a classifier that identifies general neurodivergence rather than ASD specifically. This raises the question of whether our model can truly differentiate ASD from other neurodevelopmental disorders or atypical brain patterns. Future research should explore multi-class classification models that can distinguish ASD not only from typical controls but also from other neurodevelopmental conditions.
Secondly, while our model achieved high accuracy and identified consistent biomarkers, the dataset used was limited in the amount of information given, particularly in terms of ethnic representation. This limitation potentially affects the generalisability of our findings to the broader ASD population and its potential to perpetuate healthcare disparities [75]. Future studies should prioritise the use of explicitly diverse and representative datasets to ensure that the identified biomarkers and classification methods are valid across different demographic groups.
An important future direction would be to develop a model that can highlight important regions for ASD classification on a per-MRI scan basis, rather than studying the model’s overall behavior. This approach could provide more personalised insights into the neurobiological markers of ASD for each individual, potentially revealing how these markers vary across the spectrum and among different demographic groups. This could potentially also aid clinicians with early and accurate diagnosis of ASD.
Furthermore, a multimodal approach holds significant promise. Integrating behavioural, cognitive, genetic, and structural imaging data alongside functional connectivity could provide a more holistic and nuanced picture of ASD [76]. This could lead to the identification of ASD subtypes or personalised profiles, enabling tailored interventions.
Finally, collaborative research models are needed to bridge the gap between AI development and clinical practice. Interdisciplinary teams of neuroscientists, AI researchers, and clinicians are crucial to design studies with translational impact, ensuring that research questions are clinically relevant and AI models are designed to address the complexities of real-world ASD diagnosis and care.